OpenStudy (anonymous):

Solve for x: 2 / 5x + 2 = 10

OpenStudy (anonymous):

2 / 5x + 2 = 10 now what is the 1st step to take

OpenStudy (mayankdevnani):

2/5x=8 x=8*5/2

OpenStudy (anonymous):

@mayankdevnani cant you see i want to teach her

OpenStudy (mayankdevnani):

can you solve this?

OpenStudy (mayankdevnani):

@surdawi i am also teaching....

OpenStudy (mayankdevnani):

@Stephhaniiee can you solve this?

OpenStudy (mayankdevnani):